OpenStudy (anonymous):

x/2=5/15 am I correct to say x=10/15?

OpenStudy (anonymous):

yes x=10/15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but you need to simplify

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

got ya. thanks

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

can you explain how you got 2/5?

OpenStudy (anonymous):

yes the way you set up the equation was x/2=10/15 then cross multiply you get 15x=10, then divide both sides by 15 to get the x on the left alone. and you should have x=10/15 and simplify to 2/3

OpenStudy (anonymous):

not 2/5 sorry